OpenStudy (anonymous):

What are the solutions of the system? y = x^2-3x+2 y = 4x-4 A. (1, 0) and (–6, –28) B. (1, 20) and (6, 0) C. (1, 0) and (6, 20) D. no solution A, B, C, or D? Please explain :)

OpenStudy (anonymous):

solve \[ x^2-3x+2 = 4x-4\]

OpenStudy (anonymous):

since y is expressed in two different ways you can set them equal to each other to find one or more x values that satisfies the equation x^2-3x+2 = 4x-4 x^2-3x+2-4x-4=0 x^2-7x-2 = 0 It does not seem to be easily factored so you will need the quadratic equation to solve this. Since it will likely come out to a decimal, the solution is no solution.

OpenStudy (anonymous):

Thank you @seitys. Are you positive that there is not a solution?

OpenStudy (anonymous):

there is are x values that satisfy both equations. you have to plug it into the quadratic equation to get the answer. Usually with small values <10, if it is not factor-able, it comes out to a decimal number which is not represented in your solution set.

OpenStudy (anonymous):

Okay then, thank you for your input and answer.